OpenStudy (anonymous):

Let f(x) = sin 2x.Then the derivatives of f(x) equals ?

OpenStudy (anonymous):

2cos(2x)

OpenStudy (anonymous):

hmm there is no answer in the choice.

OpenStudy (anonymous):

Choice given in the attachment.

OpenStudy (anonymous):

spetzz is right, no matter what the choices are

OpenStudy (anonymous):

Hmm but the choice given is a little confusing and complicated.I know that answer is correct.but there is no choice in it.

OpenStudy (anonymous):

i cannot open the attachment for some reason, but let me make a guess

OpenStudy (anonymous):

the answer is 2 cos(2x) so find the "double angle" formula for cos(2x) and see if that is one of the answers

OpenStudy (anonymous):

problem is there are three to choose form \[cos(2x)=1-2\sin^2(x)=2\cos^2(x)-1=\cos^2(x)-\sin^2(x)\]

OpenStudy (anonymous):

hmm the choices given got give limit wait i type out.

OpenStudy (anonymous):

forget it it is choice d

OpenStudy (anonymous):

they are not asking for the answer. just what the definition is

OpenStudy (anonymous):

wait a sec tha tis not right

OpenStudy (anonymous):

what is not right? There are x and h in the choice.That making me confuse.

OpenStudy (anonymous):

it is choice c and it was typed by someone whose math skills are limited

OpenStudy (anonymous):

You mean the answer is choice c?

OpenStudy (anonymous):

it should be \[\lim_{h\rightarrow 0}\frac{\sin(2(x+h))-\sin(2x)}{h}\] in other works you should always write \[\sin(x)\] not \[\sinx\] but whatever. your math teacher wants c.

OpenStudy (anonymous):

yes c. ask your math teacher why he or she does not think it necessary to put parentheses about the input of a function. we write f(x) not fx and sin(x) not sinx